OpenStudy (anonymous):

sqrt{(2*3^2) * 2^1/2}/ 3^-2 ? answer is 54 but how do you get it?

OpenStudy (anonymous):

like this?: \[\frac{\sqrt{2*3^{2}*2^{\frac{1}{2}}}}{3^{-2}}\]

OpenStudy (anonymous):

only 2* 3^2 is under the square root

OpenStudy (anonymous):

or maybe like this: \[\frac{\sqrt{2*3^{2}}*2^{\frac{1}{2}}}{3^{-2}}\]

OpenStudy (anonymous):

ah yes yes. ok first lets change that 2^(1/2): \[2^{\frac{1}{2}} = \sqrt{2}\] Now we can stick it inside the square root with all the other numbers: \[\sqrt{2*3^{2}}*\sqrt{2} = \sqrt{2*3^{2}*2} = \sqrt{2^{2}*3^{2}}\]

OpenStudy (anonymous):

this can be simplified a little further: \[\sqrt{2^{2}*3^{2}} = \sqrt{2^{2}}*\sqrt{3^{2}} = 2 * 3 = 6\] so the top of our fraction turns out to be 6

OpenStudy (anonymous):

so over all we have: \[\frac{6}{3^{-2}} = 6*3^{2} = 6*9 = 54\]

OpenStudy (anonymous):

i got 9 on the top but now i will look over what you did so i know how its done. thank you so much, i don''t know how to write math like that!

OpenStudy (anonymous):

if there is anything that isnt clear just let me know :)